Lawrence Industries' most recent annual dividend was $1.68 per share (d=$1.68) and the firm's required return is 11%. Find the market value of Lawrences' shares when dividends are expected to grow at 10% annually for 3 years, followed by a 7% constant annual growth rate in years 4 to infinity. a. The market value of Lawrence's shares is
Answer #1
Step-1: Present Value of dividend of Year 1 to Year 3
Year Dividend Discount Factor (1.11^-n) Present Value Dividend (working)
                1 $         1.85                                            0.901 $                    1.66 1.68*(1+.10) = $       1.85
                2 $         2.04                                            0.812 $                    1.65 1.85*(1+.10) = $       2.04
                3 $         2.24                                            0.731 $                    1.64 2.04*(1+.10) = $       2.24
Total $                    4.96
Step-2: Present value of dividend of 3 Year onwards
= (D3*(1+g)/(ke-g))*(1+Ke)^-3 Where,
= (2.24*(1+.07)/(.11-.07))*(1.11^-3) g = After 3 year Growth Rate
= $       43.81 Ke = Required return
D3 = Dividend of Year 3
Step-3: Market Value = Step-1 + Step-2
= 4.96 + 43.81
= $                 48.77
Thus, market value of share is $ 48.77